A son is 3 years younger than his sister, who, in turn is 27 years younger than their mom. In 11 years their mom will be 3 times the age of the daughter.
How old the son at the present moment.

Answer by ankor@dixie-net.com(22740)

let a = son's present age
let b = sister's
let m = mom's
:
write an equation for each statement
:
A son is 3 years younger than his sister,
a = b - 3
who, in turn is 27 years younger than their mom.
b = m - 27
In 11 years their mom will be 3 times the age of the daughter.
m + 11 = 3(b + 11)
m + 11 = 3b + 33
m = 3b + 33 - 11
m = 3b + 22
replace b with (m-27) from the 2nd statement
m = 3(m-27) + 22
m = 3m - 81 + 22
m = 3m - 59
59 = 3m - m
59 = 2m
m = 59/2
m = 29.5, not an integer so this problem is probably not valid
:
but plunging ahead anyway
b = 29.5 - 27
b = 2.5
therefore, son's age
a = 2.5 - 3
a = -.5, obviously a major flaw in the this problem
:
How old the son at the present moment. Will be born in 6 months
:
maybe the method will help you in future problems like this. CK